A body of M.I. 3 kg m2 rotating with an angular velocity 2 rad/s has the same K.E. as a mass of 12 kg moving with a velocity of
(a) 1 m/s (b) 2 m/s
(c) 4 m/s (d) 8 m/s

Grade:12

1 Answers

Vikas TU
3 years ago
Rotation of the body = 3 kg-m² (Given)
Angular velocity of the body = 3 rad/s (Given)
Weight of the mass = 27kg (Given)
Using the equation of motion -
1/2Iω² = 1/2mv²
Therefore, 1/2.3kgm². 3²rad/s = 1/2.27kgv²
= 1/2.3kgm². 9rad/s = 1/2.27kgv²
Thus v² = 1